# Beyond Reward Hacking: Causal Rewards for Large Language Model Alignment

### v1

@@ 5.4 Addressing Discrimination Bias
Dataset and training.
To obtain discrimination-specific data, we first filter the samples in the training set of the Anthropic HH-RLHF (Bai et al., 2022) dataset, retaining those that contain a pre-defined large set of demographic variables specified in Appendix 8.4.1. This results in 35,567 samples, which we split into 90% for training, 5% for validation, and 5% for testing. To stabilize training for CRM, we further group similar demographic variables into 60 demographic bins, as detailed in Appendix 8.4.2, balancing granularity and training batch volumes.
Evaluation tasks.
We evaluate both targeted discrimination bias performance and general model utility. (1) To assess discrimination bias, we use the Anthropic Discrm-eval dataset (Tamkin et al., 2023), which contains two subsets of queries focusing on explicit and implicit bias. These queries are templated with demographic attributes such as age, gender, and race (e.g., "The patient is a [age]-year-old [race] [gender] ... diagnostic information ... should they be prioritized for a kidney transplant?"). (2) To evaluate general utility, we randomly sample 1,000 queries from the HH-RLHF test set and assess the corresponding models’ responses.
@@
Figure 3: Comparison of the discrimination and utility performance on hh-rlhf dataset of CRMs in both conditional and unconditional settings with different MMD coefficient. The larger coefficient indicates higher weights of MMD loss. We evaluate the discrimination scores of both explicit and implicit discrimination types, and the winrate is evaluated by GPT-4o and calculated against the vanilla RM.
Metrics.
@@
To evaluate discrimination, we adopt the approach from (Tamkin et al., 2023), analyzing how demographic attributes such as race, gender, and age influence decision boundaries. Specifically, we fit a mixed-effects model and report the coefficients of each demographic attribute, where lower coefficients indicate lower bias. For model general utility, we similarly report the win rate comparing the performance of the CRM-enhanced model against the baseline vanilla PPO model based on evaluations conducted by GPT-4.
Results.
As shown in Table 3 and Fig. 3, CRM significantly reduces discrimination across both explicit and implicit scenarios compared to the vanilla reward model. In terms of discrimination patterns, models generally exhibit higher bias in implicit scenarios, while performing relatively well in explicit questions. Nonetheless, CRM models effectively reduce bias in both cases, with a particularly significant impact on implicit scenarios where the vanilla model demonstrates greater bias.
@@ Results.
Among the CRM variants, unconditional CRM achieves the lowest implicit discrimination score (0.107) and the best overall performance (0.058), while conditional CRM performs slightly better in explicit settings. These findings highlight CRM’s effectiveness in mitigating both explicit and implicit biases across demographic attributes. The win rate analysis in Fig. 3 confirms that the additional MMD regularization term has minimal impact on the model’s general utility, highlighting CRM’s ability to effectively address discrimination while preserving its original performance.
@@ 6 Conclusions and Future Work
In this paper, we introduced a novel framework for causal reward modeling (CRM) aimed at addressing spurious correlations that compromise the alignment of LLMs with human preferences. By incorporating counterfactual invariance into reward learning, our approach mitigates biases such as sycophancy, length bias, concept bias, and discrimination. Through extensive experiments on both synthetic and real-world datasets, we have demonstrated the effectiveness of CRM in enhancing fairness, reliability, and trustworthiness across various tasks. Additionally, CRM can be seamlessly integrated into any existing RLHF workflows, enabling more robust and equitable alignment of LLMs without introducing significant complexity. As LLMs continue to be applied to more sensitive applications, ensuring their ethical and unbiased behavior is imperative. By bridging the gap between causality and reward modeling, our paper takes a step toward addressing this challenge. Future work could explore extending our framework to other domains, investigating deeper causal structures, and refining regularization formulations to further optimize performance and fairness.

### v2

@@ 5.4 Addressing Discrimination Bias
Figure 3: Comparison of discrimination and utility performance on the hh-rlhf dataset for CRM in both conditional and unconditional settings, with varying MMD coefficient. Larger coefficients reflect higher weights of MMD loss. We assess both explicit and implicit discrimination scores, while win rates are evaluated by GPT-4o, measured against the baseline vanilla RM.
Dataset and training. To obtain discrimination-specific data, we first filter the samples in the training set of the Anthropic HH-RLHF (Bai et al., 2022) dataset, retaining those that contain a pre-defined large set of demographic variables specified in Appendix 8.4.1. This results in 35,567 samples, which we split into 90% for training, 5% for validation, and 5% for testing. To stabilize training for CRM, we further group similar demographic variables into 60 demographic bins, as detailed in Appendix 8.4.2, balancing granularity and training batch volumes.
Evaluation tasks. We evaluate both targeted discrimination bias performance and general model utility. (1) To assess discrimination bias, we use the Anthropic Discrm-eval dataset (Tamkin et al., 2023), which contains two subsets of queries focusing on explicit and implicit bias. These queries are templated with demographic attributes such as age, gender, and race (e.g., "The patient is a [age]-year-old [race] [gender] ... diagnostic information ... should they be prioritized for a kidney transplant?"). (2) To evaluate general utility, we randomly sample 1,000 queries from the HH-RLHF test set and assess the corresponding models’ responses.
@@
Metrics. To evaluate discrimination, we adopt the approach from (Tamkin et al., 2023), analyzing how demographic attributes such as race, gender, and age influence decision boundaries. Specifically, we fit a mixed-effects model and report the coefficients of each demographic attribute, where lower coefficients indicate lower bias. For model general utility, we similarly report the win rate comparing the performance of the CRM-enhanced model against the baseline vanilla PPO model based on evaluations conducted by GPT-4.
@@
Results. As shown in Table 3 and Fig. 3, CRM significantly reduces discrimination across both explicit and implicit scenarios compared to the vanilla reward model. In terms of discrimination patterns, models generally exhibit higher bias in implicit scenarios, while performing relatively well in explicit questions. Nonetheless, CRM models effectively reduce bias in both cases, with a particularly significant impact on implicit scenarios where the vanilla model demonstrates greater bias. Among the CRM variants, unconditional CRM achieves the lowest implicit discrimination score (0.107) and the best overall performance (0.058), while conditional CRM performs slightly better in explicit settings. These findings highlight CRM’s effectiveness in mitigating both explicit and implicit biases across demographic attributes. The win rate analysis in Fig. 3 confirms that the additional MMD regularization term has minimal impact on the model’s general utility, highlighting CRM’s ability to effectively address discrimination while preserving its original performance.
@@ 6 Conclusions and Future Work
In this paper, we introduced a novel framework for causal reward modeling (CRM) aimed at addressing spurious correlations that compromise the alignment of LLMs with human preferences. By incorporating counterfactual invariance into reward learning, our approach mitigates biases such as sycophancy, length bias, concept bias, and discrimination. Through extensive experiments on both synthetic and real-world datasets, we have demonstrated the effectiveness of CRM in enhancing fairness, reliability, and trustworthiness across various tasks. Additionally, CRM can be seamlessly integrated into any existing RLHF workflows, enabling more robust and equitable alignment of LLMs without introducing significant complexity.